How many grams of carbon dioxide (

CO2) are formed from 0.547 g

0.547 g of octane?


1
Expert's answer
2021-09-28T01:04:34-0400

2 C8H18 + 25 O2 → 16 CO2 + 18 H2O.

molar mass of Octane = 114.23

0.547/114.23 = 0.0048moles

Mole ratio of Octane :CO2 = 2:16

= 0.0048× 8 = 0.038moles

Molar mass of CO2 = 44.01

0.038×44.01 = 1.69g